OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u,

OpenStudy (anonymous):

unit vector in the direction of u \[=\frac{ vector ~u }{ \left| u \right| }=\frac{ -4 i-3 j }{ \sqrt{\left( -4 \right)^2+\left( -3 \right)^2} }\]

OpenStudy (anonymous):

Thank you!